OpenStudy (anonymous):

Find the exponential function f(x)=Ca^x whose graph goes through the points (0,4) and (3,32)

OpenStudy (mikey):

f(x)=4(2^x)

OpenStudy (anonymous):

whats a and whats c

OpenStudy (turingtest):

just apply the given info\[f(x)=Ca^x\]\[f(0)=Ca^0=C=4\]\[f(3)=4a^3=32\to a^3=8\to a=2\]

OpenStudy (mikey):

Take a look at Turing Test's post.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can you take a look at two more questions pls

OpenStudy (mikey):

Ok

OpenStudy (anonymous):

The graph of the function f(x)=6^(x−8) can be obtained from the graph of g(x)=6^(x) by shifting the graph of g(x) to the right 8 units The range of the function f(x) is (A,∞) what is the value of A?

OpenStudy (mikey):

So, you're asking what the minimum value for the range is?

OpenStudy (anonymous):

they want the value of A

OpenStudy (mikey):

I'm assuming that's whats being asked. So in 6^(x-8)...if x equals 8 then the exponent equals zero, right?

OpenStudy (mikey):

And as x gets larger than 8, then the exponent becomes postive and the equation goes to infinity exponentially.

OpenStudy (mikey):

So what happens to the equation when x is smaller than 8? It becomes a negative exponent. And what happens to an exponential function with a negative exponent?

OpenStudy (anonymous):

hey i got to go. if you type out the solutions i will be back to see it. thanks for the help

OpenStudy (mikey):

The graph behaves exactly like the exponential function 6^x expect that it is shifted to the right by 8 and so the y intercept is at x=8. Think about what happens to an exponential function as the exponent becomes negative. What value does it approach, but never actually gets to?